The Acmeville Metropolitan Bus Service currently charges $0.88$0.88 for an all-day ticket, and has an average of 513513 riders a day. The bus company is not earning a profit, but according to their contract with the city, they cannot cut the number of buses on the road. They must therefore find a way to increase revenues.

The bus company is considering increasing the ticket price to $1.11.1. The marketing department's studies indicate this price increase would reduce usage to 249249 riders per day.

Calculate the absolute value of the price elasticity of demand for bus tickets using the simple percentage change method. Round your answer to one decimal place.

price elasticity of demand=price elasticity of demand=

Which of the statements is true?

Demand is elastic, so decreasing ticket prices will increase revenue.

Demand is inelastic, so decreasing ticket prices will increase revenue.

Demand is inelastic, so increasing ticket prices will increase revenue.

Demand is elastic, so increasing ticket prices will increase revenue.
Answer #1

Price elasticity =(P1/Q1) *(Change in Q/Change in P)

= (0.88/513) * [(249-513)/(1.1-0.88)]

=(0.88/513)*(-264/0.22)

=-232.32/112.86

=2.05

=2.1

Since the price elasticity of demand is elastic so the company should decrease the price to increase revenue.

Option(A)